There is an important relationship between the standard normal distribution and confidence levels.  Demonstrate your understanding of this relationship by identifying the z score associated with a 94% level of confidence.  This question does not require you to use the information about the 1,496 respondents provided above.
Answer #1

Solution :-

Relationship between the standard normal distribution and confidence levels.

For a confidence interval with level C, the value = (1-C).

A 94% confidence interval for the standard normal distribution, then, the interval is (-1.881, 1.881), since 94% of the area under the curve falls within this interval.

At 94 %, Confidence level ' z ' is,

= 1 - 94 %

= 1 - 0.94

= 0.06

/2 = 0.06 / 2 = 0.03

By using Standard Normal ' z ' table ,

Z /2 = 1.881